OpenStudy (commonlizer):

What is the solution to the following equation? 5(2x − 12) + 21 = 2x + 41 5 8 10 72

OpenStudy (commonlizer):

I am taken it is 10 from what I solved

OpenStudy (anonymous):

then plug in 10 to check

OpenStudy (anonymous):

\[5(2(10)-12)+21 \overset{?}= 2(10) +41\]

OpenStudy (commonlizer):

I am some what confused

OpenStudy (commonlizer):

math is not my friend

OpenStudy (anonymous):

If x=10 solves your equation, then it will be true when you plug in 10 everywhere you see an x.

OpenStudy (anonymous):

first try the left hand side L.H.S. \[5(2(10)-12)+21=5*8+21=61\]

OpenStudy (anonymous):

and then the Right Hand Side... \[R.H.S. = 2(10)+41 = ?\]

OpenStudy (anonymous):

if they are both equal, then x=10 is a solution!

OpenStudy (anonymous):

Let me know what makes sense, and what doesn't

OpenStudy (commonlizer):

Yes 10 is the solution thank you @ Redcan

OpenStudy (anonymous):

np